A 50 kg body is applied with a force for 8 sec. After that the force is removed but the body continous in the motion for 5 sec and cover a displacement of 50m. Find the force.
1
Expert's answer
2011-07-18T07:52:22-0400
Index “1” means first part of path and index “2” the second part
V is the velocity at time 8sec
Ff is a friction force

F1 = F-Ff = ma1
a1 = (F-Ff)/m
V = a1 t2
F2 = Ff = ma2
a2 = F2/m
V = a2 t2 = F2/m t2 (*)
a1 t1 = a2 t2
(F-Ff)/m t1 = Ff/m t2
F = Ff (1+t2/t1 ) (**)
A2 = Ff S2
E1 = (mV2)/2
A2 = E1
Let put here the V from (*)
Ff S2 = m 1/2 Ff2/m2 t22
so,
Ff = (2mS2)/(t22 )
(**):
F = (2mS2)/(t22 ) (1+t1/t2 ) = (2*50*80)/52 (1+8/5) = 832 N